# Coherent coupling of YBCO superconducting resonators and 100-nm-thick YIG films

### v1

@@ 5 Conclusions
In conclusion, we have studied the microwave response of nonconductive 100-nm-thick YIG films positioned on top of a YBCO superconducting layer. Transmission spectra were acquired at different temperatures either by using a broadband CPW line or a CPW resonator. In the former case, experimental data have been essentially reproduced using well-known formulae and physical quantities describing the spin-wave excitation spectrum of the YIG film. In the latter case, the coupling extracted from the splitting of magnon-photon polaritons amounts to 230 MHz. The evolution of the polaritonic spectrum at different temperatures can be attributed to the effects of the penetration depth of the YBCO superconductor. The comparison with the results obtained with a thicker YIG film suggests a non-trivial role of both system geometry and material parameters in determining the collective coupling strength.

### v2

@@ 6 Conclusions
In conclusion, we have studied the microwave response of nonconductive 104-nm-thick YIG films positioned on top of a YBCO superconducting layer. Transmission spectra were acquired at different temperatures either by using a broadband CPW or a CPW resonator. In the former case, experimental data have been essentially reproduced using well-consolidated models and physical quantities describing the spin-wave excitation spectrum of the YIG film. In the latter case, the coupling extracted from the splitting of magnon-photon polaritons amounts to 230 MHz. The changes in the polaritonic spectrum at different temperatures can be attributed to the effects of the penetration depth in the YBCO resonator. The comparison between results obtained with YIG films having different thicknesses suggests a non-trivial role of the system geometry and material parameters in determining the evolution of the transmission spectra.